OpenStudy (mike1015):

Will medal and fan.The path of an electron, traveling at 3.7 × 104 meters/second, is making a right angle to a magnetic field. A single electron experiences a force of 2.2 × 10-15 newtons while in the field. What is the strength of the magnetic field? The value of q = -1.6 × 10-19 coulombs. 0.16 teslas 0.22 teslas 0.37 teslas 0.73 teslas

OpenStudy (michele_laino):

the magnitude \(B\) of the magnetic field, is given by the subsequent computation (Lorentz force): \[\Large B = \frac{F}{{qv}} = \frac{{2.2 \cdot {{10}^{ - 15}}}}{{1.6 \cdot {{10}^{ - 19}} \cdot 3.7 \cdot {{10}^4}}} = ...tesla\]

OpenStudy (mike1015):

thank you @Michele_Laino